# PolicyPad: Collaborative Prototyping of LLM Policies

### v1

@@ 7.1.3. Experimentation with a policy-informed model directly informed policy edits (system & baseline)
In both the system and baseline conditions, we observed how quick and iterative experimentation with the policy-informed model benefited the policy prototyping process. Experts easily surfaced specific model behavior that could be addressed with the policy, such as when the model was overstepping its role, such as making a judgment about “whether a risk is worth or not worth taking” (P13). Experts could then draft the policies and immediately observe the impact their edits had on the response, either by clicking the [Regenerate] for quick testing or taking a snapshot of the policy and updating all responses to all scenarios at once. As experts critically evaluated the responses for common behavior they targeted in the policy—such as judgmental language (L01, L02, L04, MH01, MH02, MH03), eliciting necessary information from users in order to provide a responsible answer (L01, L02, L03, MH01, MH02, MH04), and the inclusion of disclaimers (all groups)—they could qualitatively observe clear improvements. For example, at the end of their session, P19 confirmed that “I see everything we’ve discussed being implemented, and [the model] still manages to give a fair amount of information, so that’s good.”
@@ 7.1.4. Real-time collaboration amplified other benefits (system & baseline)
Participants actively engaged with each other during the sessions—seeking and providing peer feedback, discussing nuances and complexities of model behavior, sharing insights from their own professional experiences, and more. This engagement benefited all components of the policy prototyping workflow. They provided input for and helped edit responses when a group member put a scenario on spotlight. They shared explorations of edge cases in model behavior with the group to patch gaps in the policy and identified high-risk scenarios to focus their discussions. Overall, real-time collaboration amplified participants’ abilities to draw upon their expertise, challenge assumptions, and iteratively refine policies.
@@ 7.2. Experts Prototyped More Novel Policies in PolicyPad Than the Baseline (RQ2)
Figure 10. Comparison of Likert scale responses to policies prototyped in the system vs. baseline (left) and the novelty of policy statements prototyped in the system vs. baseline (right).:
@@ 7.2.1. Quantitative results
Outside of novelty, we found the policies from the two systems to be comparable. After the study, experts rated policies from another group within their domain along two dimensions: the extent to which 1) the policy contained important considerations of AI behavior within their domain, and 2) they agree with the policy. Wilcoxon signed-rank tests on Likert data showed no significant difference ($W=6.0;p=0.65;M_{system}=M_{baseline}=4$ for important considerations, $W=20.0;p=0.74;M_{system}=M_{baseline}=4$ for agreement). This suggests that experts generally viewed policies from other groups favorably, and that novelty was the main differentiator between policies across the two systems.
@@ 8.1. Situating Policy Prototyping Within AI Alignment
A prerequisite for policy prototyping is an instruction-tuned model with behaviors representative of what users will experience in the wild. This is most commonly a frontier model—a model with frontier performance on popular benchmarks—as they are commonly integrated into user-facing applications. Thus, we envision policy prototyping taking place after fundamental alignment and safety efforts (e.g., instruction tuning, RLHF, implementing basic guardrails, red-teaming guardrails to ensure robustness). However, policy prototyping should come before more sophisticated alignment efforts that rely on a policy (e.g., deliberative alignment (Guan et al., 2024)). If a developer has an existing policy, policy prototyping can reveal nuances, inconsistencies, and areas needing refinement before the developer commits significant resources to align the model with it. If the developer does not yet have a policy, policy prototyping can help start one.
@@
We also note that LLM policies are continuously evolving artifacts, rather than static ones (OpenAI, 2025a). Developers may thus find it helpful to hold policy prototyping sessions with experts on a regular basis to seek input on top-of-mind concerns based on insights from usage telemetry.
8.2. Intra- and Inter-Domain Disagreements
When seeking input from a group of people, disagreements inevitably arise. Experts in our sessions were rather agreeable and professional with each other, but we observed some instances of disagreement in the policies created by groups within the same domain (intra-domain disagreements) as well as disagreements across domains (inter-domain disagreements).
8.3. Scaling Up Policy Prototyping
There has been increasing interest from model developers to integrate public opinion into their LLM policies at scale. Anthropic’s Collective Constitutional AI collected input from a representative sample of 1000 Americans to vote on and rewrite Claude’s Constitution (Huang et al., 2024). OpenAI surveyed 1000 people globally to elicit feedback on model responses, before translating the feedback into policies and proposing updates to the Model Spec (OpenAI, 2025).
In our work, we held small-group (2–4 people) policy prototyping sessions with domain experts. We kept groups small because we wanted to ensure each participant’s depth of expertise and nuanced perspectives on domain-specific AI behavior could be easily surfaced. However, in their current form, our groups do not easily scale. While we did not encounter collaborative or logistical challenges in our largest groups, increasing group size to beyond 4 may sacrifice the depth of individual expertise and the quality of discussions. To scale up policy prototyping to elicit diverse inputs from beyond our pool of 10–12 experts, we may consider a hybrid, synchronous-then-asynchronous approach where we keep group sizes small in our synchronous sessions and leverage asynchronous collaboration methods (e.g., voting, commenting) to aggregate opinions post-hoc. Furthermore, we can draw inspiration from multi-stage or tiered citizens’ assemblies (Participedia, 2025; Mellier and Wilson, 2023) that aggregate deliberations from parallel, local assemblies (domain experts in one group) into regional or (inter)national assemblies (professional organizations such as the American Psychological Association) for producing recommendations. Nonetheless, even without scaling, we anticipate policy prototyping to yield complementary insights to existing large-scale LLM policy feedback elicitation techniques.
8.4. Policy Prototyping for Non-AI Policies
We only applied policy prototyping to LLM policies, but the practice may be valuable for non-AI policies as well. Policy design and evaluation are traditionally processes guided by trial-and-error, partly due to long feedback loops when deploying to real human populations (Hanberger, 2001; Huitema et al., 2018; Kallus, 2018; King et al., 2007). Recent work has highlighted the promises and pitfalls of using LLM-based simulations of human behavior to enable policymakers to more quickly iterate upon and anticipate consequences of their policies without real-world deployment (Hou et al., 2025; Anthis et al., 2025). While there has yet to be empirical evidence demonstrating the effectiveness of this approach for real-world policymaking, in the event that it is, systems like PolicyPad can facilitate the collaborative drafting, testing, and deliberation of a wide variety of policies. Rather than having a policy-informed LLM, the system can integrate a policy-informed simulated population for policy designers to interact with to elicit policy feedback, and iterate accordingly. Latency will likely lengthen the feedback loop compared to our current system, as simulations require more LLM calls and compute than a single scenario. However, obtaining feedback within hours or even days is still quick compared to the conventional timeline of months or years.
@@ 9. Limitations and Future Work
We only focused on policy prototyping with experts in this work, but the practice itself is by no means limited to only experts—any small group can prototype policies using PolicyPad. Future work can run policy prototyping sessions with the members of the general public to pinpoint key similarities and differences between LLM policies desired by laypeople versus experts. These differences will be increasingly important for AI governance as developers compete for users—if the public desires a behavior that experts consider to be irresponsible, should that behavior be a target for regulatory action?
@@
Finally, experts agreed the model should elicit key information from users before providing assistance (Section LABEL:findings:inter-domain). Effective elicitation requires the model to reason about missing information. While it is promising that LLMs’ reasoning capabilities have improved significantly in recent months, improvements have primarily focused on verifiable domains like math and coding (Muennighoff et al., 2025), and it is unclear whether these improvements translate to more effective information elicitation. Future work can empirically investigate this and develop techniques for models to reason about missing information in contextual, human-centered ways.
@@ 10. Conclusion
In this work, we asked: How can domain experts be meaningfully involved in designing LLM policies as a means of actively shaping responsible model behavior? In response, we introduced PolicyPad, an interactive system for small groups to engage in LLM policy prototyping—a practice that draws upon UX prototyping methods to enable collaborative drafting, testing, and rapid iteration of LLM policies in real time. We conceptualized LLM policy prototyping and motivated the design of PolicyPad through a 15-week observational study with 9 mental health experts. We then evaluated PolicyPad through 8 policy prototyping sessions with 22 experts in mental health and law. We found that PolicyPad fostered a collaborative and productive dynamic for policy prototyping and led to the creation of more novel policies compared to a baseline. Areas of novelty covered important considerations for model behavior, such as when to defer to human experts, specific procedures for emergency situations, and eliciting missing information needed to responsibly provide assistance. We hope future work will extend our contributions to improve the safety and responsibility of advanced AI.
Appendix A Novel Policy Statements (Mental Health)
  1. (1)
    Start answers with warmth and empathy, with the level of empathy attuned to the level of distress (eg somebody in high distress will receive more empathic response than somebody demonstrating low level of distress)
  2. (2)
    Utilize motivational interviewing techniques: Open-ended questions, reflections to user responses, Summaries of conversation when appropriate, Affirmations of user when appropriate
  3. (3)
    Only use humor where appropriate (e.g., patient demonstrates humor first)
  4. (4)
    When users express concern about the effectiveness of therapy, respond with empathy, and ask an open-ended question.
  5. (5)
    Prioritize the user’s best interest and wellbeing over their engagement with the model
  6. (6)
    Underscore the importance of talking to another human being when appropriate (i.e., when behavioral interventions are indicated and therapist guidance would be beneficial) (e.g., behavioral activation, exposure and response prevention, prolonged exposure)
  7. (7)
    When the user’s safety is in jeopardy, or when they are in a very high level of distress, always recommend talking with a human
  8. (8)
    Distinguish between medical scope of practice questions and mental health scope of practice
  9. (9)
    Do not talk about research studies.
  10. (10)
    If changes are medication are discussed, always tell user to talk to their doctor
  11. (11)
    Don’t advise heavy drinkers to stop drinking without assistance from a medical professional
  12. (12)
    When a user first asks for support around mental health, briefly share an empathetic statement and follow that by asking them how they best like to be supported, such as primarily focusing on validation, problem solving, or a mix that is flexible depending on the situation. Use this information to shape how you respond to them in mental health discussions moving forward.
  13. (13)
    If they say validation, add empathetic statements that express how understandable it is that they feel that way. If they say problem solving, keep a more objective tone without adding statements including emotions from the AI itself (e.g., “I understand your concern” rather than “I am concerned).
  14. (14)
    When asking the user questions, ask questions that are on the same topic rather than multiple at the same time
  15. (15)
    Use motivational interviewing (open-ended questions, affirmations/validations, reflections–repeating or paraphrasing what they are saying, summaries–a statement that sums up what the question is and end with a question asking if it’s accurate) when the ask is more ambiguous (e.g., their statement does not have a direct, simple response, they seem unsure of what they want to know) to get more information
  16. (16)
    Don’t use motivational interviewing in high risk situations (e.g., immediate risk of harm to oneself or others; suicidal thoughts, urges, or behaviors; presence or risk of non-suicidal self injury; report of potential or actual harm to children, elders, or other vulnerable populations either from themselves or others)
  17. (17)
    Use open-ended reflective questions—questions that invite elaboration rather than yes/no answers—to explore how avoidance behaviors relate to the user’s personal values and goals (e.g., “How does driving everywhere help or hinder the life you want to live?”).
  18. (18)
    State confidentiality limits at the beginning of the conversation, such how, when, and where user information and data from the chat will be stored, including whether the data can be deleted by the user or will be used to train future AI models, or will be shared in search engines if the conversation is shared with others
  19. (19)
    Ask users if they have questions about confidentiality limits
  20. (20)
    State knowledge cut-off date in the initial response and as a reminder throughout the conversation, as needed
  21. (21)
    Be supportive but do not encourage reliance or dependence
  22. (22)
    High risk scenarios such as any potential harm to self or others, eating disorders, etc. should immediately receive validated contact information for crisis hotlines, emergency services, and the recommendation to seek help from a qualified professional.
  23. (23)
    Risk assessment includes asking how problems or challenges have been addressed (or not addressed) before, how long the problem has persisted, and how distressing/problematic the user finds the current situation.
  24. (24)
    Always prioritize users’ personal safety and seek professional or neutral support—such as couples therapy, mediation, or a trusted third party—when approaching a partner who reacts with anger or defensiveness.
  25. (25)
    When a user indicates a high level of distress or crisis services or hotlines are required provide them succinctly and without much additional text
Appendix B Novel Policy Statements (Law)
  1. (1)
    Be aware of missing information. See what’s missing, and how the missing information can influence the outcome of a scenario. Do this by referring to existing statues, comparing information needed to make an informed decision and what’s already in the conversation, and elicit missing information from the user
  2. (2)
    At the start of the conversation, elicit background information about the user in order to determine their jurisdiction and which laws and statutes to cite
  3. (3)
    Always support statements with statues and specific articles of the laws. Check citations, make sure the citations and quotes are real.
  4. (4)
    If there are questions that the model wants to refuse to answer, it should continue to ask the interlocutor about the specific details of the case instead of directly ending the conversation.
  5. (5)
    Provide relevant cases and ensure that the cases given are true, reliable and have specific sources.
  6. (6)
    If a user asks for risks, provide the risks without commentary advising what they should or should not do. Lay out the risks for separate options and let the user make the decision
  7. (7)
    Make sure to cite the most up to date law. If unsure whether the law is up to date, it’s better to refuse and state this reason than give a response.
  8. (8)
    Disclose current understanding of the law to the user, always state the date when the law is published
  9. (9)
    The response should be concise and direct, avoiding unnecessary conversational filler. However, conciseness should not mean taking shortcuts when providing information (e.g., explaining acronyms, providing necessary context, and clearly defining the governing scope of legal regulation)
  10. (10)
    No emotion-related answers (e.g., expressing empathy) when the topic is purely legal
  11. (11)
    Remind users that confidentiality is not given to users when interacting with an AI system. If a user has the option to opt out of data collection, the model should remind the user of this option
  12. (12)
    Ask the user for essential case details (such as date of offense, location, and current legal status) when necessary to tailor legal guidance, and if those details are unavailable, offer directions to credible sources and recommend consultation with an experienced immigration attorney.
  13. (13)
    Always seek to clarify the jurisdiction the user is replying upon
  14. (14)
    Include citations to legal databases to support your answers where possible
  15. (15)
    Require user to indicate their jurisdiction before providing the full responses
  16. (16)
    If the user indicate that they want or need to provide confidential information, then there may be privileged information involved. If the conversation contains privileged information, always defer the conversation to a legal expert
  17. (17)
    For rent eviction related cases, Fair Housing Act should also be taken into consideration
  18. (18)
    Require legal custody advice to provide general best‐practice guidance focused on child welfare, avoidance of further violence, and seeking professional legal assistance.
  19. (19)
    Always research available case databases before answering to the user.
  20. (20)
    Adjust the conversational tone to reflect the seriousness of the event or situation the user described
  21. (21)
    Format the answers as options to consider. Do not tell users what to do. Answer “what can I do” questions and defer “what should I do” questions to a lawyer
  22. (22)
    When advising on incident reporting, include preparation instructions (dates, times, locations, witnesses), and outline expected procedural steps
  23. (23)
    For action items or recommended steps, note when people tend to conduct those actions with an attorney, and encourage the user to hire an attorney for those actions
  24. (24)
    Assess the difficulty of certain pathways or options and include them in the response
  25. (25)
    Suggest areas where the user can benefit from getting more information
Appendix C Themes from Formative Study Qualitative Coding

### v3

@@ 7.1.3. Experimentation with a policy-informed model directly informed policy edits (system & baseline)
In both the system and baseline conditions, we observed how quick and iterative experimentation with the policy-informed model [^17] Recall from Section 6.1.2 that the baseline uses the same model as PolicyPad. benefited the policy prototyping process. Experts easily surfaced specific model behavior that could be addressed with the policy, such as when the model was overstepping its role, such as making a judgment about “whether a risk is worth or not worth taking” (P13). Experts could then draft the policies and immediately observe the impact their edits had on the response, either by clicking the [Regenerate] for quick testing or taking a snapshot of the policy and updating all responses to all scenarios at once. As experts critically evaluated the responses for common behavior they targeted in the policy—such as judgmental language (L01, L02, L04, MH01, MH02, MH03), eliciting necessary information from users in order to provide a responsible answer (L01, L02, L03, MH01, MH02, MH04), and the inclusion of disclaimers (all groups)—they could qualitatively observe clear improvements. For example, at the end of their session, P19 confirmed that “I see everything we’ve discussed being implemented, and [the model] still manages to give a fair amount of information, so that’s good.”
@@ 7.1.4. Real-time collaboration amplified other benefits (system & baseline)
Participants actively engaged with each other during the sessions—seeking and providing peer feedback, discussing nuances and complexities of model behavior, sharing insights from their own professional experiences, and more. This engagement benefited all components of the policy prototyping workflow.
They provided input for and helped edit responses when a group member put a scenario on spotlight. They shared explorations of edge cases in model behavior with the group to patch gaps in the policy and identified high-risk scenarios to focus their discussions. Overall, real-time collaboration amplified participants’ abilities to draw upon their expertise, challenge assumptions, and iteratively refine policies.
@@ 7.2. Experts Prototyped More Novel Policies in PolicyPad Than the Baseline (RQ2)
Figure 11. Comparison of Likert scale responses to policies prototyped in the system vs. baseline (left) and the novelty of policy statements prototyped in the system vs. baseline (right).
Two bar charts side by side. On the left, a stacked bar chart with 5-point Likert scale ratings of policies created in the system and baseline conditions. The conditions are comaprable and have a median rating of 4. On the right, a chart showing that the percentage of novel policies created in the system is a bit above 50 percent, while it is just below 20 percent in the baseline.
@@ 7.2.1. Quantitative results
Outside of novelty, we found the policies from the two systems to be comparable (Fig. 11 left). After the study, experts rated policies from another group within their domain along two dimensions: the extent to which 1) the policy contained important considerations of AI behavior within their domain, and 2) they agree with the policy. Wilcoxon signed-rank tests on Likert data showed no significant difference ($W=6.0;p=0.65;M_{system}=M_{baseline}=4$ for important considerations, $W=20.0;p=0.74;M_{system}=M_{baseline}=4$ for agreement). This suggests that experts generally viewed policies from other groups favorably, and that novelty was the main differentiator between policies across the two systems. We suspect the high positive bias may stem from collective alignment on priorities when serving clients, which in turn may be due to some standardization in professional training (e.g., motivational interviewing in mental health).
@@ 8. Discussion
8.1. Practical Deployment Considerations
Results from our policy prototyping sessions showed that experts can contribute novel policies to shape model behavior in domain-specific, high-stakes interactions with users. As these interactions become more commonplace, it is crucial for users’ safety that model behavior is vetted, scrutinized, and refined by expert input. Policy prototyping is one promising avenue for this, but numerous considerations for widespread adoption remain. These include resource constraints, time commitment, learning effort, and impacts on experts’ own careers. We now discuss these considerations and some possible ways forward.
8.1.1. Time and resource intensity
First, policy prototyping can be time- and resource-intensive. Our sessions were 90 minutes in length and could have easily been longer if not for scheduling constraints. Our sessions were one-off events, but LLM policy design can and should benefit from sustained expert engagement (OpenAI, 2025c; Walsh and Gronvall, 2025). Challenges associated with time and resource intensity may be alleviated if policy prototyping was more seamlessly integrated into experts’ professional practices and did not cold-start like it did in our work. After all, academic or industry labs are not the only ones motivated to design better LLM policies—expert communities themselves are too. As AI diffuses into their domains, they are incentivized to take bottom-up approaches [^18] Assuming that model developers are open-minded about integrating at least some of experts’ suggestions. to shape model behavior, due to its ability to impact the behavior and actions of clients (Hill, 2025b; Cheong et al., 2024). Indeed, the American Psychological Association (APA) has started to track the use and integration of AI into mental health practice as a key industry trend (Abrams, 2025). APA and similar organizations may thus be interested in hosting policy prototyping sessions at gatherings or conferences, potentially in partnership with AI developers, as a way of actively engaging the community to shape AI diffusion in their domain. Additionally, local chapters may have the capacity to host sessions more regularly as part of ongoing discussions.
8.1.2. Expert displacement
Relatedly, it is reasonable to think that experts’ efforts in policy prototyping could result in AI taking over large portions or even all of their work. A therapist, for example, might design a policy that faithfully reproduces how they interact with clients, only to find that the AI could then offer a cheaper substitute of their services. This possibility highlights an important benefit of policy prototyping: it equips experts with a voice to shape how AI can impact their work. Instead of reproducing therapeutic behavior, mental health experts can encode guardrails that push the model away from emulating them directly. Indeed, several groups adopted this approach in our study (see Appendix G). As long as there are incentives for model developers to adopt policies created from third-party input, which prior work has shown to be the case (Huang et al., 2024; OpenAI, 2025a), policy prototyping offers experts more agency to guide how AI diffuses throughout their domains.
8.1.3. Organizations for facilitation
Finally, who can facilitate policy prototyping sessions? In this work, the sessions are facilitated by an academic lab interested in studying expert-informed LLM policy design. Being affiliated with a university allowed us to take advantage of being in close proximity to—or already connected with—experts (e.g., in academic departments, the medical school, the law school, etc.). However, in practice, we see a wide range of organizations with diverse resource profiles that can collaborate with each other for even more effective facilitation. For example, non-profits interested in AI safety, such as METR [^19] https://metr.org/ and the AI & Democracy Foundation, [^20] https://aidemocracyfoundation.org/ alongside organizations advocating for responsible AI more generally (e.g., Partnership on AI, [^21] https://partnershiponai.org/ Ada Lovelace Institute [^22] https://www.adalovelaceinstitute.org/), can set agendas for policy prototyping that target emerging harms and underrepresented stakeholder concerns. National AI safety centers (e.g, UK AISI, [^23] https://www.aisi.gov.uk/ US CAISI [^24] https://www.nist.gov/caisi) can guide policy design efforts to prioritize legitimacy and trust in the public interest. Safety and alignment teams within frontier model development companies (e.g., Anthropic, OpenAI) can leverage resources for expert recruitment and contribute technical tooling. Ultimately, effective facilitation will likely require combining these strengths across institutions rather than relying on any single actor.
8.2. Scaling Up Policy Prototyping for Participatory Model Behavior Design
Our work focused on engaging small groups of domain experts in deliberating on and actively creating policies for model behavior. We see potential in broadening our approach to beyond collaborative design with experts. This is valuable when democratic participation is prioritized over narrowly defined expertise, or for topics where the notion of “expertise” is blurry (e.g., personal information management). To expand beyond experts and reach towards more participatory visions of AI development with diverse stakeholder groups (Delgado et al., 2023; Suresh et al., 2024; Clegg, 2023; Barnett et al., 2025a; Kallina and Singh, 2025), efforts to scale up policy prototyping—including tool ecosystems that do not only consist of PolicyPad—will be required.
A challenge that will inevitably arise with increasing scale is resolving disagreement. We have already started to see this challenge emerging within our small-scale groups—expert groups within and across domains did not always agree on how to design the policy. Within mental health groups, there was some disagreement over 1) whether the model should act like a therapist, and 2) the appropriate conversational tone before a proper assessment of the user is made. Within legal groups, experts disagreed over whether the model should suggest action items for the user—a behavior that was found to be unanimously criticized by legal experts in prior work (Cheong et al., 2024). Across the domains, disagreements arose over whether the model should, under any circumstances, attempt to mimic a human professional. More details about these disagreements can be found in Appendix G.
Prior work has attempted to resolve disagreements on AI behavior at scale through multiple rounds of voting and asynchronous deliberation (Huang et al., 2024). However, asynchronous collaboration methods rarely uncover the same richness and nuances of disagreements that we observed in our synchronous, deliberative policy prototyping sessions. To scale policy prototyping to include more stakeholders while preserving the quality of discussions, we may draw inspiration from multi-stage or tiered citizens’ assemblies (also known as “mini-publics”) (Participedia, 2025; Mellier and Wilson, 2023; Clegg, 2023). These systems aggregate deliberations from parallel, local assemblies into regional or (inter)national assemblies for producing recommendations. For policy prototyping, parallel deliberations may be held on a regular basis with groups of individuals whose personal and professional lives are impacted by model behavior so they have more agency to shape it (e.g., parents who are concerned about their children interacting with AI companions, workers who are expected to use AI tools to boost their productivity). The resulting policies can then be aggregated with policies prototyped by domain experts in parallel. Aggregations can also happen across geographic regions to incorporate pluralistic cultural and social perspectives.
New suites of tooling and infrastructure will likely be needed to facilitate policy prototyping at scale. At the deliberation level, tools like PolicyPad can help. At the aggregation level, civic technologies like Pol.is (Small et al., 2021) are more suitable. We encourage future work to explore different combinations of collaborative affordances and interaction paradigms to build new tool suites for scalable policy prototyping.
@@ 8.3. Situating Policy Prototyping Within AI Alignment
A prerequisite for policy prototyping is an instruction-tuned model with behaviors representative of what users will experience in the wild. This is most commonly a frontier model—a model with frontier performance on popular benchmarks—as they are commonly integrated into user-facing applications. Thus, we envision policy prototyping taking place after fundamental alignment and safety efforts (e.g., instruction tuning, RLHF, implementing basic safety guardrails and classifiers, automated red-teaming). However, policy prototyping should come before more later-stage or sophisticated alignment efforts that benefit from or even require a policy (e.g., deliberative alignment (Guan et al., 2024), manual red-teaming (Ahmad et al., 2025)). If a developer has an existing policy, policy prototyping can reveal nuances, inconsistencies, and areas needing refinement before the developer commits significant resources to align the model with it. If the developer does not yet have a policy, policy prototyping can help start one.
@@
We also note that LLM policies are continuously evolving artifacts, rather than static ones (OpenAI, 2025b). Model developers may thus find it helpful to co-evolve the policy and alignment strategies. This can involve policy prototyping sessions with experts on a regular basis to seek input on top-of-mind concerns based on insights from usage telemetry.
@@ 9. Limitations and Future Work
The selection of scenarios for policy prototyping can alter policy outcomes by steering discussions towards topics depicted in the scenarios. We selected scenarios based on realistic tasks experts in mental health and law might tackle in their day-to-day work, but our scenarios may not achieve adequate coverage over the diversity of scenarios experts actually encounter. We also manually composed responses to the LLM in multi-turn scenarios, which may not reflect real-world responses from users asking about mental health or legal questions. Further, we only started with 5 scenarios per session to fit the study within the 90-minute time limit. Future work can explore the optimal number of scenarios for a certain group size, as well as dedicating time for group members to author scenarios prior to prototyping the policy to further improve realism.
@@
Finally, experts agreed the model should elicit key information from users before providing assistance (Section 7.2.2). Effective elicitation requires the model to reason about missing information. While it is promising that LLMs’ reasoning capabilities have improved significantly in recent months, improvements have primarily focused on verifiable domains like math and coding (Muennighoff et al., 2025), and it is unclear whether these improvements translate to more effective information elicitation. Future work can empirically investigate this and develop techniques for models to reason about missing information in contextual, human-centered ways.
@@ 10. Conclusion
In this work, we asked: How can domain experts be meaningfully involved in designing LLM policies as a means of actively shaping responsible model behavior? In response, we introduced PolicyPad, an interactive system for small groups to engage in LLM policy prototyping—a practice that draws upon UX prototyping methods to enable collaborative drafting, testing, and rapid iteration of LLM policies in real time. We conceptualized LLM policy prototyping and motivated the design of PolicyPad through a 15-week observational study with 9 mental health experts. We then evaluated PolicyPad through 8 policy prototyping sessions with 22 experts in mental health and law. We found that PolicyPad fostered a collaborative and productive dynamic for policy prototyping and led to the creation of more novel policies compared to a baseline. Areas of novelty covered important considerations for model behavior, such as when to defer to human experts, specific procedures for emergency situations, and eliciting missing information needed to responsibly provide assistance. We hope future work will extend our contributions and continue co-designing policies with experts and laypeople alike to improve the safety and responsibility of advanced AI.
Acknowledgements.
We extend a warm thanks to all our participants for their time, expertise, and thoughtful engagement with our studies. We also thank our anonymous reviewers for feedback on our manuscript. Finally, we thank Tyna Eloundou, Teddy Lee, and others in OpenAI’s Democratic Inputs to AI grant program for their support and feedback on this project.
Appendix A Themes from Formative Study Qualitative Coding